Volume 2, Number 3 July 1993

TEX and TUG NEWS

for and by the TEX community

A Publication of the TEX Users Group Electronic version TEX and TUG NEWS

Mission Statement

The TEX Users Group (TUG) provides leadership:

1. to encourage and expand the use of TEX, , and related systems

2. to ensure the integrity and portability of TEX, , and related systems

3. to foster innovation in high-quality electronic document preparation

TEX and TUG NEWS is a newsletter for TEX and L^{A}TEX users alike: a forum for exchanging information, tips and suggestions; a regular means of communicating news items to one another; a place where information about TEX and TUG can be quickly disseminated.

Throughout the newsletter “TEX” is understood to mean TEX, L^{A}TEX, AMS-TEX, and other related programs and macros. TEX and TUG NEWS is produced with the standard L^{A}TEX distribution, and is to be as portable a document as possible.

The entire contents of this newsletter are being placed in the public domain. The source file of this issue will be placed in the aston, shsu, and stuttgart archives, as well as at the heidelberg, labrea, and ymir archives. Copying and reprinting are encouraged; however, an acknowledgement specifying TEX and TUG NEWS as the source would be appreciated.

Submissions to TEX and TUG NEWS should be short, the macros must work, and the files must run without special font or graphics requirements: this is to be a portable newsletter (the new font selection scheme has not yet been implemented). Correspondence may be sent via e-mail to tug@math.ams.org with the subject line

NEWSLETTER.

TEX Users Group  Phone: +1 (805) 963-1338
P.O. Box 869 FAX: +1 (805) 963-8358
Santa Barbara, CA 93102 USA E-mail: tug@math.ams.org

TEX is a trademark of the American Mathematical Society.

TEX and TUG NEWS (ISSN 1065-240X) is published quarterly by the TEX Users Group, 735 State Street, P.O. Box 869, Santa Barbara, CA 93102, U.S.A. The 1993 dues for individual members are $60.00 for regular members and $30.00 for students.

Second-class postage paid at Santa Barbara, CA, and additional mailing offices. Postmaster: Send address changes to TEX and TUG NEWS, TEX Users Group, P.O. Box 869, Santa Barbara, CA 93102, U.S.A.

Membership in the TEX Users Group is for the calendar year, and includes all issues of TUGboat and TEX and TUG NEWS for the year in which membership begins or is renewed. Individual membership is open only to named individuals, and carries with it such rights and responsibilities as voting in the annual election.

Editorial

News! News! And still more News! Maybe it’s a sign of the excitement which the Annual Meeting generates, but there has been an avalanche of news releases and announcements crossing my screen these days. Things I’ve had no room for include:

• the Textures 1.6 release

• Arbortext and Solaris 2.x for TEX users

• Bernard Gaulle has announced a new release of french.sty v.3.20

• Timothy van Zandt has a new release of PSTricks out, available also from

TUG

• there’s a new thing called ACROBAT — if we’re lucky, we’ll be seeing something about it at Aston

• and here’s a last-minute arrival: a July release of something called TEXhelp: The On-Line TEX Handbook, from Arvind Borde

Question: should we be mentioning new releases in TTN? What do you think? People who cruise the newsgroups read about them all the time: it seems a reasonable thing to bring some of those announcements here to TTN. What sorts of things do you want to know about?

We have a list of the conference presentations as they are currently known, and we have a list of the courses and instructors for the week before, and the week after the conference itself. Lots is happening, that’s for sure! And remember to bring some discs if you want to copy some of the updates/upgrades to software from the fabulous Aston Archive.

See you at Aston. And for those who won’t be able to make it, the Proceedings issue of TUGboat will bring the conference to you in a few months’ time.

Christina Thiele

Editor, TEX and TUG NEWS

Oct. issue  Sept. 15th
Jan. issue (1994) Dec. 1st
Apr. issue  March 15th

Typographer’s Inn

Well-versed

The late Brian O’Nolan (Flann O’Brien), writing his satirical column Cruiskeen Lawn (as Myles na gCopaleen) in the Irish Times in the 1930s commented:

Having considered the matter in—of course—all its aspects, I have decided that there is no excuse for poetry. Poetry gives no adequate return in money, is expensive to print by reason of the waste of space occasioned by its form, and nearly always promulgates illusory concepts of life. [My italics]

Several times I have been asked to help with the formatting of poetry. There are several ‘poem.sty’ files around, but none of them seem to tackle the fundamental problem of coping with overlong lines. The poet, untrammelled by having to consider such mundane factors as typography, quite rightly expects the printer to make as decent a shot as possible at reproducing the poem, but all too often the ‘printer’ is nowadays an inexperienced DTP operator, with scant knowledge of typographic practice.

It is particularly evident when setting in short measures, but the only satisfactory solution I have been able to come up with is the following kind of layout (this is a verse of a hymn which had to be set in 10/11 pt to 12\frac{1}{2} pc):

6 And in the garden secretly, And on the cross on high, Should teach his brethren, and To suffer and to die. [inspire

What I haven’t been able to find is the way to get TEX to do this by itself. The usual pint of Guinness at Aston for the first solution. I should mention that when there is insufficient space at the end of the next line, and if there is sufficient space at the end of the preceding line, then the excess word should be brought up to the preceding line! The alternate indenting is handled by the code:

\newcount\plines
\everypar={\global\advance\plines by1
           \ifodd\plines\else\quad\fi}
\obeylines%

so this extra space also needs to be taken into account.

Curiously, as I was in the middle of writing this, a question about verse formatting was posted to comp.text.tex, so it would be nice to come up with some answers.

When is an em not an em?

Talking of comp.text.tex, there has been a debate recently about the size of an em. A user wanted to know how big an em was, and if that size was related to the current point size of type. Once the confusion was settled over the separate meaning of a pica em (always 12 pt, regardless of what type you are setting in: this is TEX’s ‘pc’ unit of measurement), it became obvious that there is substantial disagreement, or at least misunderstanding, over the term ‘em’.^{1}

Historically, an em was the width of a capital ‘M’ (approximately, anyway). As more and more type designs deviated from this, it became necessary to fix the value at some standard: normally 1 em is the same as the current pointsize: that is, if you are setting in 8 pt type, then the instruction ‘leave 1 em space’ means ‘leave 8 pt space’. Anyone who has ever set metal type knows this from the fact that an em quad is a square-section slug of typemetal of side x pt, where x is the typesize in points.

However, does not obey this convention: if you process the following code, you will find that the width of what claims is 1 em in cmbx10 is actually 56.19455pt.

\newbox\embox \newdimen\emwidth
\setbox\embox=\hbox{\normalsize\hbox to1em{\hfill}}
\emwidth=\wd\embox
\showthe\emwidth

I’m not sure why it does this: I have never found a use for it, and have always had to be explicit when writing macros to make sure that if the typographer’s spec says ‘Chapter Number in 8/9 pt Times Bold, run in 2 em to Chapter Title in 8/9 pt Times Roman’ that the result leave exactly 16 pt between CN and CT, otherwise the client will be upset.

Can anyone shed some more light on this? Perhaps it’s some devious and hidden difference between typographical practice in North America and in Englishspeaking parts of Europe.

Fonts

I’ve just managed to get a real PostScript laser printer near me: up until now I’ve been making do with Ghostscript for drafts and trusting to a well-tried local typesetter for repro pulls and camera-ready. I still think the default interletter spacing on the Adobe fonts is too tight. I don’t have a copy of ATF Garamond, but if anyone out there can lay their hands on a specimen book, and measure an alphabet, I’d be glad of the opportunity to compare it with the one sold by Adobe.

1And there we have a classical example of where to put the punctuation. Presumably the MLA would prefer me to finish the sentence with ‘em.’

New Publications

Peter Schmitt

                                         schmitt@awirap.bitnet
                           a8131dal@awiuni11.edvz.univie.ac.at

Helmut Kopka and Patrick W. Daly. A Guide to L^{A}TEX: Document Preparation for Beginners and Advanced Users. Addison-Wesley (UK) Ltd., 1993. 448pp. US$34.50. ISBN 0-201-56889-6.

• This book is based on two well-known German books by Helmut Kopka: L^{A}TEX — Eine Einführung, and L^{A}TEX — Erweiterungsmöglichkeiten.

Francesc Rosselló Llompart. L’ABC del TEX.

Palma de Mallorca: UIB-DMI Departament de Ciències Matemàtiques i Informàtica, Universitat de les Illes Balears, 1991.

Reinhard Wonneberger. L^{A}TEX. 3., durchgesehene und erweiterte Auflage. (Addison-Wesley Kompaktführer) Bonn etc.: Addison-Wesley, 1993. XVI, 166pp. DM 29,80. ISBN 3-89319-589-0.

• This is the third (expanded) edition of a concise German L^{A}TEX reference (previous editions 1987 and 1988). The main addition is a section on german.sty, which is widely used in German-speaking countries.

Electronic publications

Hubert Partl, Elis. Schlegl and Irene Hyna. Una Descripción de L^{A}TEX. • Spanish translation by Tomas Bautista of well-known German introduction to L^{A}TEX, L^{A}TEX-Kurzbeschreibung. Includes material on the spanish style option. Available in Spain from ftp.eunet.es (also known as goya.

uu.es) in the directory info/unix/text/TeX/spanish/tutorial, as well

as from the main TEX archives.

Latest issues of other TEX newsletters

Les Cahiers GUTenberg: No. 14: EuroTEX’92: Proceedings of the 7th European TEX Conference. Prague, Sept. 14–18, 1992. Pp. vi + 330. No. 15, avril 1993, 61pp. ISSN 1140-9304.

GUST: Zyszyt 1 (1993). [March issue 35 + 7; A4 format.] First issue of the newsletter GUST from the Polish user group Grupa Użytkowników Systemu TEX. Editors are: Włodek Bzyl and Tomasz Przechlewski. Address is: Instytut Matematyki UG, Wita Stwosza 56,

80-952 Gdańsk. E-mail: matwb@halina.univ.gda.pl.

MAPs: issue 93.1, 1993, 233pp. [Dutch-speaking user group, NTG publication].

Feature Article

International quotations _{Johannes Braams PTT Research J.L.Braams@research.ptt.nl}

Introduction

At the conference in Portland last year, Christina Thiele asked me if I was willing to write an article for TTN about quotation marks. She said that she suspected that there would be many people out there who would want to know how to produce quotation marks in a little piece of, say, French text that they include in their document that is otherwise written in English. All this is of course covered in the Babel language-specific style files, but for those who don’t want to use Babel but do need the occasional quotation mark I wrote this article.

Let me start with a disclaimer: I am not familiar with all the typographic conventions that are in use in the various countries all over the world. Also I can only give you macros for those languages for which a language-specific Babel file exists. So, there may be more conventions. If you know of them, please inform me, so that I can enlarge my collection.

Yet another disclaimer: If you have access to fonts with the DC/EC (or Cork) encoding, you do not need most of the hackery in this article. Those fonts include the necessary characters, so they need not be constructed.

The first part of this article^{1} deals with the macros needed to typeset the quotation marks needed for various languages. First, I’ll introduce some internal hackery, then the macros that really produce quotation marks. I also give some shortcuts to make the macros easy to use. Finally, I will give some examples of the use of the macros. In the examples, I will show the source text as well as the typeset result.

Some help macros

Because the macros in this part of the code are not intended for use in a document, we’d like to ‘hide’ them from the user. Usually this is done by changing the category code of a character with category code ‘other’ to a letter. More often than not the ‘@’ is chosen for this purpose. Sometimes the underscore is used for this purpose as well. Here I too will use the ‘ ’. Of course the category code needs to be reset later so we store the current value in \uscatcode:

       \chardef\uscatcode=\catcode`\_
       \catcode`\_=11\relax

1When the source of this article is run through L^{A}TEX it produces a file called quoting.tex which contains all the code that is described and used in the article.

In some languages we need to lower quotation marks to the baseline. For this purpose we use the macro \set_low_box. It has one argument and uses \box0 to gives us its result. This macro comes from german.tex.

It firsts typesets a comma in box register 2 and its argument in box register 1. Then it computes the distance that the argument has to be lowered to reach the baseline. Finally, it lowers the contents of box register 0 — using box register 0 again for the result — and adjusts the values for the height and depth of the box:

    \def\set_low_box#1{\setbox2\hbox{,}\setbox0\hbox{#1}%
      \dimen0\ht0 \advance\dimen0 -\ht2%
      \setbox0\hbox{\lower\dimen0 \box0}\ht0\ht2 \dp0\dp2}

Using macros for quotation marks sometimes disturbs the spacefactor, so we also need a macro to preserve it. This macro also stems from german.tex. It first checks if it is executed in horizontal mode; if that is the case, the current spacefactor is stored in the macro \_SF. Outside horizontal mode the macro \_SF is empty. Then \save_sf_q typesets its argument and resets the spacefactor:

    \def\save_sf_q#1{{\ifhmode
      \edef\_SF{\spacefactor\the\spacefactor}\else
      \let\_SF\empty \fi \leavevmode #1\_SF}}

Producing the quotation marks

In languages such as Dutch, German and Czech, the opening quotes are traditionally typeset at the baseline:

    \def\loq{\protect\_loq}
    \def\_loq{\save_sf_q{\set_low_box{''}\box0\kern-.04em}}

In German, the closing quotes are also different from what is provided by TEX. They use something that looks like the Eglish opening quotes as closing quotes. Obviously, if one didn’t do anything about it, the spacing would be wrong. Therefore we need yet another macro, \icq:

    \def\icq{\protect\_icq}
    \def\_icq{\save_sf_q{\kern-.07em``\kern.07em}}

In French typography, a very different kind of quote mark is used, the socalled ‘guillemets’. To realise these guillemets, various macros have been floating around the net. According to the french.sty package by Bernard Gaulle, one of the oldest definitions is:

    \def\oog{\protect\_oog}
    \def\_oog{\leavevmode\ifdim\lastskip>0pt\unskip
               \penalty-9\hskip0.35em minus 0.35em\fi

               \raise .27ex\hbox{$\scriptscriptstyle\ll$}%
               $\,$\nobreak\ignorespaces}
    \def\ocg{\protect\_ocg}
    \def\_ocg{\leavevmode\ifdim\lastskip>0pt\unskip
               \penalty10000\fi
               \nobreak$\,$\leavevmode
               \raise .27ex\hbox{$\scriptscriptstyle\gg$}}

But this does not really give the result that the French would like. Therefore, if you have access to the L^{A}TEX symbol fonts it is better to use the following definition for the guillemets:

    \chardef\lg='050
    \chardef\rg='051
    \def\og{\protect\_og}
    \def\_og{\hbox{\ly\lg\kern-0.2em\lg\kern+0.2em}}
    \def\cg{\protect\_cg}
    \def\_cg{\hbox{\ly\kern+0.2em\rg\kern-0.2em\rg}}

In the code above the control sequence \ly is used. It is an internal macro from the old font selection scheme in L^{A}TEX. When you use the NFSS you will have to define it this way:

    \ifx\undefined\selectfont
    \else
      \def\ly{\fontfamily{lasy}\fontseries{m}
              \fontshape{n}\selectfont}
    \fi

Easy usage

In the previous section a couple of macros have been defined to make it possible to use various quotes. But it would be nice if one didn’t have to do so much typing each time they were used. To provide easy access, it is common use in language-specific files to introduce active characters. For the macros presented here, we could introduce three active characters, the ", < and the >:

    \def\dq{"}\catcode`\"=\active
    \def\lt{<}\catcode`\<=\active
    \def\gt{>}\catcode`\>=\active

As you may have noticed, I saved a copy of the non-active version of each character in a control sequence. These are needed later on, when the active character has inspected its argument and decides that it needs to insert the non-active version of itself.

We use the active " to access the low opening quotes and the German closing quotes; the other two are used to produce the guillemets. Here is the definition of the active characters:

    \def"#1{\ifx#1`\loq{}\else
            \ifx#1'\icq{}\else\dq#1
            \fi\fi}
    \def<#1{\ifx#1<\ifmmode\lt\lt\else\og{}\fi\else\lt#1\fi}
    \def>#1{\ifx#1>\ifmmode\gt\gt\else\cg{}\fi\else\gt#1\fi}

But, be careful when introducing new active characters. You have to make sure that they get deactivated at the right moment. Therefore, we need to add them to macros such as \dospecials and — in case you use L^{A}TEX — \@sanitize. A safe way of doing this was found by Bernd Raichle. It involves using an extra macro \add_special:

    \chardef\atcatcode=\catcode`\@
    \catcode`\@=11\relax
    \def\add_special#1{\begingroup
        \def\do{\noexpand\do\noexpand}
        \def\@makeother{\noexpand\@makeother\noexpand}
        \edef\x{\endgroup
          \def\noexpand\dospecials{\dospecials\do#1}
          \expandafter\ifx\csname @sanitize\endcsname\relax \else
            \def\noexpand\@sanitize{\@sanitize\@makeother#1}
          \fi}
        \x}
    \catcode`\@=\atcatcode\relax

Once that macro is defined we use it to tell TEX to treat our active characters with caution:

    \add_special\"
    \add_special\<
    \add_special\>

Wrapping up

These macros were defined while the category code of the ‘ ’ was changed. It must not be forgotten to undo that change:

    \catcode`\_=\uscatcode\relax

Examples

A quotation in Dutch might look like this:

   Hij zei: "`Ga je mee?''.  Hij zei: ”Ga je mee?”.

But in French (using the definition that uses the L^{A}TEX symbol font) it would be:

   Il disait: << Tu va? >>.  Il disait: (( Tu va? )).

Whereas in Italian it might look like:

   Lui dice: >>Andiamo?<<.  Lui dice: ))Andiamo? (( .

To show the difference between the two definitions given for the guillemets, this is what the old version looks like:

   Il disait: \oog\ Tu va? \ocg.  Il disait: ^{≪} Tu va?^{≫}.

“Hey — it works!”

Jeremy Gibbons University of Auckland

                                            jeremy@cs.aukuni.ac.nz

Welcome to “Hey — it works!” This column is designed for little tricks — even ones that seem trivial or inelegant — that others may find useful. If you have any such items, please send them in to me at the address above.

Paul Hafner, in the Department of Mathematics and Statistics here at Auckland, pointed out a problem with Hans Guesgen’s macros in the last issue for interspersing headings in bibliographies. Hans was working from an older version of L^{A}TEX, and an update to L^{A}TEX in January 1992 broke Hans’ macro; the fix is to add the definition

\def\@bibitem#1{\item\if@filesw \immediate\write\@auxout
       {\string\bibcite{#1}{}}\fi\ignorespaces}
to bibheadings.sty.

The first piece in this issue is a brief note from Klaus Lagally about Peter Schmitt’s margin mark macros in TTN 1,3. Next we have an item by Claudio Beccari on fine adjustments to the spacing of alignments, expanding on his own comment in TTN 2,1. We conclude with a short piece from Charles Wells on giving arbitrary labels to equations.

More on margin marks _{Klaus Lagally University of Stuttgart lagally@informatik.uni-stuttgart.de}

Here are a few comments on “Marking lines in the margin”, TTN 1,3 p. 12 and TTN 1,4 p. 10, by Peter Schmitt.

The macros as given will inhibit hyphenation. This can be corrected by redefining:

\long\def\readword#1 {#1\strut\putmark\space\next\relax}

If the margin mark is a \vrule of sufficient height and depth the \strut is no longer necessary, not even at the beginning of a paragraph, as TEX will now be in horizontal mode anyway.

How to redefine \putmark to get the mark into the right margin is left as an exercise to the interested reader.

Correct spacing for tables and arrays _{Claudio Beccari Politecnico di Torino beccari@polito.it}

Every L^{A}TEX user knows how to insert vertical spacing in tables, arrays and equation arrays by means of the optional argument of the \\ command. I suppose that most users do not like to explore the inner depths of L^{A}TEX, so that it is possible that some users do not realize that the \\ command behaves quite differently in equation arrays compared with tables and arrays.

In the eqnarray environment (with or without asterisk) the \\ command with an optional argument — a dimension given in square brackets — actually inserts that vertical space in addition to the default row spacing defined by the format file (in fact this default spacing is given by the \jot dimension, which

is defined in lplain.tex, not in book, report, article, or letter, nor in the

size-related subsidiary files); of course, anybody can assign this \jot a different value, but most users simply don’t do it.

The \\ command optional argument behaves in a different way in the tabular and array environments; this optional argument is used to increment the depth of the strut that implicitly is inserted in the first cell of every row and in the only cell produced by means of the \multicolumn command. Therefore it is impossible or difficult to control vertical spacing after every \hline and \cline command, and after those rows that contain at least one cell whose depth is greater than the implicit strut depth.

With the set of L^{A}TEX or plain TEX predefined commands, in the first case there is no workaround, except, perhaps, inserting a whole row of empty cells, which may result in too much white space. In the second case, it is possible to proceed empirically until one finds the right amount of extra space to insert as the optional argument to the \\ command. In any case, if you try to modify the implicit strut dimensions by redefining \arraystretch, you end up with a table or an array that contains too much white space.

Both cases, on the other hand, may be dealt with very nicely if some new struts are defined and inserted in the right places (in addition to the implicit strut) in order to control the vertical space very finely. In my tables and arrays I generally use the following two struts, but in special cases I might define several similar invisible objects:

\newcommand\T{\rule{0pt}{2.6ex}}              % Top strut
\newcommand\B{\rule[-1.2ex]{0pt}{0pt}}        % Bottom strut

The ‘top strut’ ensures that its row is at least 2.6ex high; the ‘bottom strut’ ensures that it is at least 1.2ex deep.

Compare the following two ruled tables; you will notice in the second one that horizontal rules are not too close to capital letters and that fractions (even if they are set in textstyle and are therefore smaller than usual) do not require guesswork in order to keep them at a reasonable distance from the following rule:

Laplace transforms

f(t)  (p) no struts
δ(t) 1 optional arg.
cos ω_{0}t \frac{p}{p^{2}+ω_{0}^{2}}

Laplace transforms

f(t)  (p)  with struts
δ(t)  1  optional arg.
cos ω_{0}t \frac{p}{p^{2}+ω_{0}^{2}}

Notice that a bottom strut in the second table is inserted in the denominator of the fraction, not to the side of the fraction, otherwise its extra space would be masked by the fraction depth.

The second table is obtained with the following commands:

\newcommand\T{\rule{0pt}{2.6ex}}              % Top strut
\newcommand\B{\rule[-1.2ex]{0pt}{0pt}}        % Bottom strut
\begin{tabular}[t]{|c|c|l|}
   \hline
   \multicolumn{3}{|c|}{Laplace transforms\T\B}\\\hline
   $f(t)$       \T & $(p)$      \B & with struts  \\\hline
   $\delta(t)$     & $1$        \T & optional arg.\\[.5ex]
   $\cos\omega_0t$ & $\frac{p}{p^2+\omega_0^2\B}$ &    \\\hline
\end{tabular}

The \\ command optional argument is used after the third row, because it does not contain oversized cells; otherwise, it would be better to control vertical spacing with struts only.

Instead of locally defining the top and bottom struts, as done in the example, it is possible to put the definitions in the preamble, or, even better, to define a command that defines them, so that the definition may be kept local to a particular environment or group:

\newcommand\setTBstruts{\def\T{\rule{0pt}{2.6ex}}%
                        \def\B{\rule[-1.2ex]{0pt}{0pt}}}

so that it is possible to get them to work with a simple call to \setTBstruts only where it is needed.

Labelling equations by names _{Charles Wells Case Western Reserve University cfw2@po.cwru.edu}

Sometimes you want to give one or two of the main results of your paper a name. If the result is in the form of an equation, you might want it to be labeled by the name you gave it instead of by a number. The following four lines define two new commands, \labeq and \elabeq, that produce a labeled equation for which you choose the label. You can also refer to it later by its label or its page number. The equation does not use up an equation number: if the equation before the labeled equation is number n, the equation after the labeled equation will be number n + 1.

\makeatletter
\def\labeq#1{\def\@currentlabel{#1}\label{#1}$$}
\def\elabeq{\eqno{\mbox{(\@currentlabel})}$$}
\makeatother

You would use it like this:

\labeq{MAIN}
x^2+y^2=z^2
\elabeq

which would print as

x^{2} + y^{2} = z^{2}  (MAIN)

You could refer to it by writing

See equation~(\ref{MAIN}) on page~\pageref{MAIN}.

This would print as “See equation (MAIN) on page 12.” Note that you must put the parentheses in yourself.

This definition is a genuine hack since it depends on internal L^{A}TEX code that is not guaranteed to be the same in future revisions.

(L^{A})TEX News

Announcing OzTEX 1.5 _{Andrew K. Trevorrow Australian National University akt150@huxley.anu.edu.au}

OzTEX 1.5 is now available at the following FTP sites, as well as from TUG:

       rsphysse.anu.edu.au in pub/oztex/newoztex
       ftp.tex.ac.uk  in pub/archive/systems/mac/oztex

It should also be at midway.uchicago.edu. There have been a lot of changes in version 1.5; here are the most significant ones:

• OzTEX now has a “big” TEX (32-bit halfwords). The values of mem_max,

     mem_top, pool_size and font_mem_size can all exceed 64K. Other TEX

parameter limits have also been increased.

• Multiple folders for TEX input files and .tfm files are allowed.

• OzTEX supports all types of printers (not just PostScript printers).

• Non-resident PostScript font files can be downloaded on demand.

• OzTEX supports background printing and typesetting.

• New keyboard short cuts make it easier to proofread a .dvi file. (E.g., you can hit the space bar to scroll forwards through a document.)

• The “OzTEX User Guide” has been updated and expanded.

• OzTeX is now shareware. More info about the shareware fee is in OzTEX’s Help menu; or contact the author for more details. All earlier versions remain freeware.

BibTEX HyperCard stack _{Evan Antworth Summer Institute of Linguistics evan.antworth@sil.org}

HyperBibTEX is a Macintosh HyperCard application for managing bibliography databases in a format compatible with BibTEX (as defined in Lamport’s L^{A}TEX book and Patashnik’s document “BibTEXing”). A HyperBibTEX bibliography stack can be created either by manually making new cards or by importing an existing .bib file. When you want to make a bibliography for a document, you simply export selected cards from the stack to a .bib file and process it with BibTEX. HyperBibTEX has facilities for searching, sorting, marking cards and data validation.

HyperBibTEX requires HyperCard version 2; version 2.1 is recommended because of its superior error handling. It should run fine under System 6, but the Balloon Help facility requires System 7. It also runs under HyperCard Player.

HyperBibTEX version 0.9.7, released on 3 May 1993, is an upgrade that fixes various bugs (including a serious bug in version 0.9.5) and adds a number of new improvements and features. If you are presently using an older version of HyperBibTEX, I strongly recommend that you upgrade to version 0.9.7. HyperBibTEX is freeware and is available by anonymous FTP from:

mac.archive.umich.edu:

     /mac/hypercard/organization/hyperbibtex0.97.sit.hqx

sumex-aim.stanford.edu:

     /info-mac/card/hyper-bib-tex-097.hqx

midway.uchicago.edu

     /pub/OzTeX/bibtex/HyperBibTeX097.sea.hqx

Or you can order HyperBibTEX on diskette from:

International Academic Bookstore e-mail: academic.books@sil.org
7500 W. Camp Wisdom Road  phone: 214/709-2404
Dallas, TX 75236  FAX: 214/709-2433

U.S.A.

Send all correspondence about HyperBibTEX to Evan Antworth, Academic Computing Department at the Summer Institute of Linguistics (phone 214/709-3346).

UK TEX archive on CD-ROM Rich Morin

Prime Time Freeware (PTF) 2-1 contains a (300 MB, compressed) snapshot of the UK TEX Archive, as of late December, 1992. Issue 2-2 will contain a snapshot of files changed between that time and late May, 1993.

PTF is a semi-annual collection of UNIX-related freeware source code and documentation. Each issue consists of two ISO-9660 CD-ROMs, bound into a 50+ page booklet. PTF is particularly useful for programmers who do not have FTP access, but many programmers use PTF to save disk space and avoid annoying FTP searches and retrievals.

PTF 2-1 contains 1,200 MB of compressed archives, unpacking to more than 3,000 MB of source code and documentation, current as of January 15, 1993. The top five items, in order of size (KB), include: the UK TEX Archive (299,508), SRC Modula-3 (82,198), NetLib Archive (math and sim.) (81,650), ICOT (5th Gen. AI Code) (78,792), and Scheme Language (36,540).

Inquiries and orders may be sent (no e-mail orders, please) to:

Prime Time Freeware 370 Altair Way, Suite 150 Tel: +1 408-433-9662
ptf@cfcl.com  Sunnyvale, CA 94086 USA Fax: +1 408-432-6149

The CTAN archives _{George D. Greenwade Chair, TWG on TEX Archive Guidelines bed gdg@SHSU.edu}

The TUG Technical Working Group on TEX Archive Guidelines (TWG-TAG) was established with the mandate of “develop[ing] guidelines for the effective management and utilization of major TEX archives, and to initiate communication among the maintainers of the existing archives for the purpose of coordination and synchronization.”

The result of this effort has been the creation and development of the Comprehensive TEX Archive Network (CTAN). In cooperation with DANTE’s archive at Stuttgart and the UK TEX Users Group’s archive at Aston, Sam Houston State University (Huntsville, Texas, USA) joins in what is very possibly the largest coordinated archive of file offerings available for network-based retrieval. The goal of the CTAN is to have available comprehensive archives of TEX-related files (sources, macros, utilities, drivers, pre-packaged complete sets for given operating systems, etc.) within parallel directory and filename structures accessible by common user interfaces. Within a very brief time lag, the holdings at one CTAN host should be precisely the same as on any other CTAN host. Moreover, the interface, command syntax and available local commands to assist users should be the same regardless of which host you are connected to.

Each CTAN host updates its collection with files and packages submitted for archiving, and places them into the pre-established directories within the CTAN structure. From there, they are passed along to the other CTAN hosts as they mirror one another on a daily basis. Additionally, each of the CTAN hosts may serve as a point of entry for contributions directly uploaded to it, which will then be passed along to the other hosts in the network.

As noted in the preliminary announcement for the upcoming TUG meeting, the CTAN is a topic which will receive wide coverage, as well as demonstrations for those interested. Presently, the Internet File Transfer Protocol (FTP) interfaces for anonymous use are parallel between CTAN hosts; eventually electronic mail interfaces to the CTAN archives will also be parallel (and reported as soon as feasible). For users with FTP client access, you are welcome to view and use any of the established hosts:

Host Name  Base CTAN Directory

             ftp.uni-stuttgart.de  /pub/tex/
             ftp.tex.ac.uk         /pub/archive/
             ftp.shsu.edu          /tex-archive/

via anonymous FTP (login as “anonymous” and use your complete electronic mail address as your password). Beyond these base directories, the holdings should parallel one another very closely.

You Rang …?

Robert Becker

Department of Mathematics, Room 2–332 Massachusetts Institute of Technology

77 Massachusetts Avenue Cambridge, MA 02139–4307

FAX: (617)253–4358

Welcome to the second column of You Rang …? As stated in the previous issue of TTN, this column will attempt to satisfy two major needs: first, hardware/software tips and tricks for the general user; second, tips on how to install versions of TEX and their support packages (dvips, for example). Questions come from you, the reader, by postal mail (not e-mail or COD) or by FAX (see the above number). I definitely want to hear from you. Tell me what you’re trying to do and, space and interest providing, a solution will be published.^{1} I cannot emphasize enough how important your input is to the TEX Users Group. I’m looking forward to seeing your questions and comments!

How can I get an indent after a sectioning command?

Sectioning commands all use a construction called \@startsection, which is defined in latex.tex. Let’s take a look at it (I’ve numbered the lines for explanatory purposes):

1.\def\@startsection#1#2#3#4#5#6{\if@noskipsec \leavevmode \fi
2.  \par \@tempskipa #4\relax
3.  \@afterindenttrue
4.  \ifdim \@tempskipa <\z@ \@tempskipa -\@tempskipa \@afterindentfalse\fi
5.  \if@nobreak \everypar{}\else
6.    \addpenalty{\@secpenalty}\addvspace{\@tempskipa}\fi \@ifstar
7.    {\@ssect{#3}{#4}{#5}{#6}}{\@dblarg{\@sect{#1}{#2}{#3}{#4}{#5}{#6}}}}

On line 4 at the end, see the \@afterindentfalse command. That says that after the section header is typeset, no indent is allowed in the first paragraph. To allow the indent, copy the above command to a separate file (say, commands.sty) in the same directory as the document on which you are work-

ing. Now change \@afterindentfalse to \@afterindenttrue. Last, make sure

you call commands.sty after all other style files are called, and — voilà!

            \documentstyle[11pt,leqno,commands]{report}

1I will attempt to answer your question, even if it doesn’t get published. If I can’t answer it, you’ll get a pointer to more information.

How do I get an abstract in the right place in L^{A}TEX using

\maketitle?

This is actually a trick question. Using the \maketitle command depends on having a titlepage environment. In addition, the titlepage environment only exists in the article and report document styles. Here’s the general structure (taken from Lamport, page 164^{2}):

\begin{titlepage} % Start of Titlepage area
  \title{Title of article or report}
  \author{Author 1\thanks{Acknowledgement of support 1.}
          Company of Author 1\\
          Address of Author 1                          \and
          Author 2\thanks{Acknowledgement of support 2.}
          Company of Author 2\\
          Address of Author 2                          \and
          Author 3\thanks{Acknowledgement of support 3.}
          Company of Author 3\\
          Address of Author 3
         }
  \date{Date of Paper}
\end{titlepage}
\maketitle                   % Set the Titlepage
\begin{abstract}             % Abstract comes after the Titlepage
   Text of abstract.
\end{abstract}

Let’s tackle this step by step.

\begin{titlepage} This starts the environment. \title{Title} This does the obvious. In a long title, you can use the \\

command to tell L^{A}TEX where to begin a new line.

 \author{Author(s)} Obvious.

\and When there are multiple authors, this command goes between each (you don’t have to — it’s a matter of style: you could simply type

     \author{Author 1, ..., and Author n}). In addition, if each author

has an address you wish to specify, typeset it as in the above example, with \\ to start new lines where appropriate. In the above example, each author will be centered with the corresponding address and will look something like this (only larger because they’re authors):

2Leslie Lamport: L^{A}TEX: A Document Preparation System, User’s Guide & Reference Manual, Addison-Wesley, ISBN 0-201-15790-X.

Author 1 Address of Author 1

Author 2 Address of Author 2

Author 3 Address of Author 3

\thanks{} This command puts important information in a footnote at the bottom of the titlepage. Note that the footnote uses symbols (∗, †, ‡ …, ‡‡), not numbers, to differentiate them from footnotes in the main text.

\date{} This command puts a date you specify into the titlepage and is optional. If you leave it out, L^{A}TEX will put in today’s date. You can specify no date by typing \date{}.

\end{titlepage} Ends the titlepage environment.

\maketitle This command tells L^{A}TEX to put together the titlepage. If this isn’t in the document, the titlepage doesn’t get typeset. Remember to put it after the titlepage environment: otherwise the abstract goes before the titlepage on a separate page!

\begin…\end{abstract} The abstract goes here. It will be typeset in a

smaller typeface than normal text. The titlepage style. This is only necessary in the article document style

as the report document style automatically pulls titlepage.sty in (and book and letter document styles don’t use titlepage). This option specifies that the titlepage and abstract go on separate pages.

How do I change the line spacing?

L^{A}TEX has the construction \baselinestretch which is used with \renewcommand in the following manner:

1. In the preamble, use \renewcommand{\baselinestretch}{num} where

num is a whole number (−1 works, but anything less has the same effect).

2. After the preamble, the above command only works after a typesize change,

so you need to use something like this:

    \renewcommand{\baselinestretch}{num}\large\normalsize

Plain TEX has the length \baselineskip and is used as follows:

         \baselineskip=num_{1}plusnum_{2} minus num_{3}

where num_{1} is the amount of skip you want (12pt skip with 10pt text), num_{2} is the amount it can stretch (3 or 4pt), and num_{3} is the amount it can shrink (1 or 2pt). These plus and minus options are necessary so TEX can break pages without overfull or underfull messages appearing. Another general query is double-spacing; while a 24pt baselineskip is technically “double-spaced”, the appearance can be much improved by using a skip of only 80% over single spacing. In a case of 10pt text, try a skip of 20pt and see if that’s acceptable.

TUG ’93 Updates

TUG ’93 program (July 26–31) Aston University, Birmingham UK

Monday, 26 July

10.00–12.30: Tutorials

Either • Introduction to L^{A}TEX (what it is and what it is not) or • Flavours of TEX: a brief tour

Either • Getting TEX: how to set up and maintain a TEX system for yourself and your friends or • Fonts for TEX: how fonts are accessed by TEX, and the many possibilities for obtaining ‘typeset output’

2.00–6.00: Orientation/Keynotes

• A new typesetting system: is it really necessary?

• Brum for beginners

• TEX from \indent to \par

• The future of TEX and TUG

Speakers: Peter Abbott, Bogusław Jackowski, Joachim Lammarsch, Marek Ryćko, Christina Thiele.

Workshops: L^{A}TEX3 or Virtual fonts

Tuesday, 27 July

8.45–12.30: Multilingual

• A format compilation framework for European languages

• Language-dependent ligatures

• Russian TEX issues

• Typesetting Catalan texts with TEX

• Working group report

Speakers: Gabriel Valiente Feruglio, Robert Fuster, Yannis Haralambous, Irina Makhovaya, John Plaice, Larry Siebenmann

2.00–6.00: Tools/Margins

• An abstract model for tables

• Bibliography prettyprinting and syntax checking

• LexiTEX: context-sensitive citations

• Maps in

• Mixing TEX and SGML: a recipe for disaster?

Speakers: Nelson Beebe, Frank Bennett, Peter Flynn, Daniel Taupin, Xinxin Wang, Derick Wood

Workshops: MakeIndex or Problem

Wednesday, 28 July

8.45–1.30: Futures

• A future for TEX

• Beginner’s guide to DSSL

• Building a future for TEX^{1}

• Building a TEX-based multi-window environment adapted to research work

• Document centred applications with GRIF

• NTS: the future of TEX?

Speakers: Martin Bryan, Roger Hunter, Michel Lavaud, Jean Paoli, Philip Taylor, Michael Vulis

Thursday, 29 July

8.45–1.05: Fonts, Maths

• A PostScript font installer written in TEX

• A versatile TEX device driver

• On the readability of maths typesetting

• Scalable outline fonts

• The Khmer script tamed by the Lion (of TEX)

• The spacing around mathematics

• Virtual fonts in a production environment

Speakers: Michael Doob, Yannis Haralambous, Berthold Horn, Alan Jeffrey, Minato Kawaguti, David Murphy, Craig Platt, Larry Siebenmann

2.00–6.30: TUG/Archives

• TEX Users Group general meeting

• The comprehensive TEX archive network — CTAN

Speakers: George Greenwade, etc.

Workshops: BibTEX or Multilingual

Friday, 30 July

8.45–12.30: Macros/Didot

• Electronic type design

• Fundamental TEX macros for processing structured documents

• News from MusicTEX • Sorting in BLUe

1Text in slant font implies an unconfirmed or tentative title.

• Syntactic sugar

• Teaching typography — the Didot project

Speakers: Mike Daines, Mary Dyson, Jonathon Fine, Kees van der Laan, Daniel Taupin

2.00–3.30: Didot

• Document design

• Wrap-up

Speakers: Richard Southall, etc.

TUG ’93 courses (20 July–6 August) _{Carol Hewlett TUG ’93 Committee}

The following courses will be held the week before the TUG ’93 Annual Meeting, and the week after the meeting.

1. Beginning/Intermediate TEX: 20–24 July Tutor: Michael Doob

2. Intensive L^{A}TEX: 31 July–3 August and 2–5 August (2 courses) Tutor: Malcolm Clark (31 July–3 August)

3. Advanced TEX and macros: 20–24 July Tutor: Chris Rowley

4. TEX output routines: 2–4 August Tutor: Philip Taylor

5. Modifying L^{A}TEX 2.09 style files: 31 July–1 August Tutor: Sue Brooks

6. logos: 21–23 July Tutor: Eric-Jan Vens

7. fonts: 2–6 August Tutor: Yannis Haralambous

8. Using the TEX family for setting maths: 31 July Tutor: Rosemary Bailey

9. Book design in TEX: 22–23 July Tutor: Philip Taylor

10. Doing more with L^{A}TEX 2.09: 24–25 July Tutors: Michel Goossens and Frank Mittelbach

11. Beyond Computer Modern: using other fonts in TEX: 22–23 July Tutor: Yannis Haralambous

12. TEX and PostScript: 22–23 July Tutor: Sebastian Rahtz All courses will be held at Aston University, Birmingham, UK. For further details, please contact:

Carol Hewlett  E-mail: hewlett@lse.ac.uk
Computer Service, LSE  Phone: +44 71-955 7926
London, WC2A 2AE, UK FAX: +44 71-955 7001

Reports on Meetings

NTG’s Lustrum Meeting _{Kees van der Laan Chairman, NTG}

On June 10th, the NTG had its first Lustrum (fifth anniversary) meeting at KNMI De Bilt. Some 60 members (about 25% of the membership) attended the meeting, which was held under beautiful weather. A message entitled “Congratulations from TEX Users Group” was read aloud on behalf of Christina Thiele, who could not attend the meeting. The NTG board underwent some changes: Johannes Braams was re-elected, and Erik Frambach joined the board as a new member, replacing Theo Jurriens, all by acclamation. Theo Jurriens was thanked for his all-too-short presence on the board, and one of his ideas was realized at this meeting: to pay special attention to the Flemish members. Philippe Vanoverbeke has accepted the task of NTG’s Belgian commissionar. This function is informal for the time being. Some formal issues were discussed during the “Other business”.

A walk through NTG’s history lane was presented by Kees van der Laan, interactively, meaning that attendees could introduce their own appreciations of the past or their visions about the future. For the most recent past, Frans Goddijn gave the speaker some rest by announcing NTG’s bulletin board, prepared and maintained by Henk de Haan and himself.^{1} My own vision on the future did not concentrate so much on the future of TEX, but more on the production and consumption of information by the community at large, and TEX’s role in this, next to other tools to be expected to develop.

After the busy lunch — there is always much to talk about, to page through, etc. — we had our meat: Huib van Krimpen, a famous Dutch typographer, and Frank Blokland, a coming man in font design. We also had our own contributions: Rein Smedinga about his experience in preparing a book via L^{A}TEX, and instead of the announced Theo Jurriens, who could not make it, Johannes Braams presented a paper he and Walter van der Laan (no relation) had presented at the TUG’92 meeting at Portland: ‘Writing reports with more than a hundred people”. Ça va sans dire, that this was an excellent program, and we all enjoyed it very much. Frank Blokland decided on the spot to become a member of NTG, because TEX, etc. intrigued him so much, and he — with his professional craftsman background — expects to learn a lot from the other side.^{2}

1A Dutch article on the issue appears in MAPS93.1, pp. 31–36.

The speakers got a big hand for their fine contributions. The president invited the attendees to extend this to the editorial team, the working groups for their fruits during the past five years, and to the board and the membership for the achievements we have made. The host was thanked for their kind hospitality, and that it was a bit overdone — to make up for those high temperatures outside. Enough heat was on already. We finished up with an informal lustrum drink, and the traditional dinner with the speakers, and members to hang on.

TUG Board Activities

1993 TEX Users Group Election _{Barbara Beeton for the Elections Committee}

The terms of all 15 members of the first elected Board of Directors will expire on December 31, 1993. The election to choose the new Board members will be held this fall, and nominations are invited.

The Bylaws provide that “Any member may be nominated for election to the Board by submitting a nomination petition in accordance with the TUG Election Procedures. Election of the directors shall be by written mail ballot of the entire membership, carried out in accordance with those same Procedures. Each director will hold office for a term of two (2) years. Directors may be re-elected for successive terms.”

The name of any member may be placed in nomination for election to this office by submission of a petition, signed by two other current (1993) members, to the TUG office at least 30 days prior to the election. A petition form follows this announcement; forms may also be obtained from the TUG office and will be available at the annual meeting, to be held this year at Aston University, Birmingham, UK.

Along with a petition form, each candidate is asked to supply a passportsize photograph, a short biography, and a statement of intent to be included with the ballot; the biography and statement of intent together may not exceed 400 words. The deadline for receipt at the TUG office of petitions and ballot information is September 1, 1993.

Some changes have been proposed in the structure of the Board and the terms of Board members. As most Board members attend the annual meeting, and making arrangements to attend meetings at other times of the year is often

2If not we from him.

Nomination for 1993 TUG Board of Directors Election

Only current (1993) TUG members are eligible to participate. The signatures of two (2) members are required in addition to that of the nominee. Type or print names clearly, exactly as they appear in the most recent TUG membership list or on a TUG mailing label; new members should enter the name which they used on their membership application form. Names that do not exactly match the TUG records will not be accepted as valid.

The undersigned TUG members propose the nomination of:

Name of nominee  (signature)  (date)

for the position of Member of the TUG Board of Directors, for a term beginning January 1, 1994.

Members supporting this nomination

Nominated by  Signature  Date

(please print)

Return this petition to the TUG office (FAXed submissions will also be accepted). Petitions must be received in the TUG office no later than September 1, 1993.

TEX Users Group FAX: 805-963-8358 Nominations for 1993 Election P.O. Box 869 Santa Barbara, CA 93102-0869 U.S.A. difficult, it has been proposed that the term be shifted so that the terms of outgoing members end, and those of new members begin, with the annual business meeting. This arrangement should assure a smooth transition, with both old and new members able to attend at the same time.

A second proposed change would alter the lengths of some terms for one time only, so that not all members of the Board are replaced at the same time, but only one-third. For Board members elected this year, it would mean that five will have terms that end with the 1995 annual meeting, five with the 1996 annual meeting, and five with the 1996 annual meeting. Thereafter, all new terms will be for three years. All positions will be elected this year on the same basis, and the members of each of the three groups chosen by lot by the disinterested person or organization whose services are procured to count the ballots and certify the results.

Subject to approval by the present Board at their annual meeting, these new procedures will be in effect for the upcoming election.

Ballots will be mailed to all members about 30 days after the close of nominations. Marked ballots must be received no more than 6 weeks following the mailing; the exact date will be noted on the ballots.

Ballots will be counted by a disinterested party not part of the TUG organization. The results of the election should be available by the end of November.

Notices

• Technical Council _{Michael Ferguson for the Technical Council}

The Technical Council would like to solicit ideas and volunteers for the creation of “Special Interest” Technical Working Groups. A current example is

WG-92-04 (SI-TWG) Title: Special Interest TWG TEX for the Disabled

Mandate: The primary purpose of this working group is as a forum for those people interested in using and/or enhancing TEX to serve the needs of the disabled.

     TWG Chair: T.V. Raman (raman@cs.cornell.edu)

Please contact Michael Ferguson, Yannis Haralambous or Alan Hoenig, at the TUG meeting in Aston or by email at

       Michael Ferguson (Chair): mike@inrs-telecom.uquebec.ca
       Yannis Haralambous:  yannis@gat.citilille.fr
       Alan Hoenig:  ajhjj@cunyvm.cuny.edu

• Conference Planning Committee_{Jackie Damrau for the Conference Planning Committee}

We are soliciting the membership for volunteers to host our annual meetings. The meeting this year in Birmingham, England is our first meeting outside of North America, and hopefully will enable many more European members to attend the annual meeting. In keeping with this approach, we would like to promote a meeting sequence in which meetings would occur in a different area of the world each year, thus giving all TUG members an opportunity to attend an annual meeting periodically without undue expense. It is our hope that members from different countries will volunteer.

What is the process for deciding where meetings will be held? The committee solicits bids from members, and submits proposals to the TUG Board, who make the final decision.

What is involved? Using the recorded experience of past conference organizers, the committee has compiled a minimal checklist for volunteers to complete and return to the committee for an initial bid to host a meeting. Along with this checklist, the committee plans to provide organizers a summary of tips — a sort of “things to do” and “things to avoid”.

Would you like to host an annual meeting at your site? If you are interested, please contact committee chair, Jackie Damrau, at

Jackie Damrau, Chair  Phone: 214-708-6048
SSC Laboratory FAX: 214-708-5143
Mailstop 1011, Ste. 125  E-mail: tugcpc-l@irlearn.bitnet

2550 Beckleymeade Avenue Dallas, Texas 75237, USA

Next Board Meeting

24–25 July, at TUG ’93

Aston University, Birmingham, England

TEX Users Group 1993 Course Schedule

Beginning/Intermediate TEX*  Intensive Course in L^{A}TEX*
Santa Barbara** Oct. 18–22  Ottawa  August 23–27
Santa Barbara  Feb. 7–11  Santa Barbara** Oct. 25–29
Santa Barbara  Jan. 31–Feb. 4
Modifying L^{A}TEX Style Files*  Adv’d TEX and Macro Writing*
Santa Barbara  Feb. 28–Mar. 4  Santa Barbara** Nov. 1–5
Santa Barbara  Feb. 14–18
TEX for Publishers  Practical SGML and TEX*
Washington, D.C. Nov. 12  Santa Barbara** Nov. 8–9
Book and Document Design  SGML and TEX for
with TEX  Publishers*
Boston Sept. 23–24  New York Nov. 10

*Lab classes — computers will be provided for all students. **NOTE: These classes, previously scheduled for Boston, have been moved to Santa Barbara.

• TUG courses are small, with 8–15 students in most classes.

• Dates and locations subject to change — contact the TEX Users Group at 805-963-1338 or send a FAX to 805-963-8358.

On-Site Courses in TEX and L^{A}TEX from TEX Users Group

• Courses in TEX, L^{A}TEX, SGML and TEX, PostScript, or TEX for Publishers tailored to the needs of your group

• Courses at every level from beginning to advanced

• Five full days of instruction at your site

• One-week course fee includes all instructor fees and expenses plus textbooks and other materials for up to 15 students

Upcoming Events

26–30 July

TUG ’93:

“World Wide Window on TEX” 14th Annual Meeting, Aston University, Birmingham, UK.

Chris Rowley, ca rowley

                            @vax.acs.open.ac.uk

Malcolm Clark

                            malcolmc@wmin.ac.uk

15 Aug.

EP94, RIDT94 deadline: Call for papers for major joint conference in Darmstadt, Germany, 11–15 April 1994.

                            EP94: ep94@gmd.de
                            RIDT94: ridt94@irisa.fr

23–24 Sept.

DANTE:

9th general meeting of DANTE, including free tutorials on various themes. Kaiserslautern.

Klaus Uttler

                            uttler@rhrk.uni-kl.de

28–29 Sept.

Electronic Books 1993: First annual meeting, at the Sheraton New York, New York City. Pre-conference workshops on 27 Sept.

                            Meckler@jvnc.net

Phone: (203) 226-6967 In USA: 1-800-632-5537 FAX: (203) 454-5840

18 Nov.

NTG 12th Meeting: “(L^{A})TEX User Environment”. Den Bosch, OCÉ.

Gerard van Nes

                            vannes@ecn.nl

Spring 1994 NTG 13th Meeting: “(L^{A})TEX, , and tools education.” Groningen, at RUG.

Gerard van Nes

                            vannes@ecn.nl

Note: Also consult the “Calendar” in the previous issue of TUGboat for more dates and details.

TEX and TUG NEWS Table of Contents

Editorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Typographer’s Inn Peter Flynn . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

New Publications Peter Schmitt . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

Feature Article:

International quotations Johannes Braams . . . . . . . . . 5

“Hey — it works!”

Jeremy Gibbons . . . . . . . . . . . . . . . . . . . . . . . . . . 9

(L^{A})TEX News Announcing OzTEX 1.5 Andrew K. Trevorrow . . . . . . . . 13 BibTEX HyperCard stack Evan Antworth . . . . . . . . . . 13 UKTEX archive on CD-ROM Rich Morin . . . . . . . . . . . 14 The CTAN archives George D. Greenwade . . . . . . . . . . 15

You Rang …?

Robert Becker . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

TUG ’93 Updates TUG ’93 program (July 26–31) . . . . . . . . . . . . . . . . . . 19 TUG ’93 courses (20 July–6 August) Carol Hewlett . . . . . 21

Reports on Meetings NTG’s Lustrum Meeting Kees van der Laan . . . . . . . . . 22

TUG Board Activities 1993 TEX Users Group Election . . . . . . . . . . . . . . . . . 23 Nomination form . . . . . . . . . . . . . . . . . . . . . . . . . 24 Notices

Technical Council . . . . . . . . . . . . . . . . . . . . . . 25 Conference Planning Committee . . . . . . . . . . . . . . 26

TEX Users Group 1993 Course Schedule . . . . . . . . . . . . . . . . . 27

Upcoming Events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

Volume 2, No. 3, 1993

Eight Great Reasons to Join the TEX Users Group

1. Publications: All members receive 4 copies of TUGboat, The Communications of the TEX Users Group; 4 copies of TEX and TUG NEWS (TTN); and a membership directory. One of the four TUGboat issues contains the proceedings of the annual meeting, so if you can’t make the meeting, you can still keep up to date with the latest developments. As well, TUG publishes special single-topic items in its TEXniques series; some are guides and summaries of TEX and L^{A}TEX; others offer descriptions of major software packages (e.g., PICTEX, EDMAC).

2. Books and Software: A “one-stop shop,” TUG carries most TEX macro packages as well as TEX publications and products. The Radel collection of public domain TEX software for microcomputers is now available through TUG, too. Members benefit from a 10% discount on all items available from TUG.

3. Membership Networks: TUG maintains and distributes a complete membership directory annually. Members are listed alphabetically, geographically and by institution. These lists facilitate easy member-to-member communication within TUG’s worldwide network.

4. Information Referrals: TUG can connect you with site coordinators — specialists in implementations of TEX on various computer architectures — as well as other resource volunteers. These experts are available to answer your questions about TEX, TEX-related software and other technical matters.

5. Annual Meetings: TUG’s Annual Meetings bring TEX users together to learn the latest in TEX applications and innovations through seminars, talks and informal gatherings. They offer an opportunity to connect with other TEX users at every level. Each conference is informative, but always informal and friendly — the perfect forum for valuable interaction and exchange of ideas on TEX.

6. TEX Worldwide: When you belong to TUG you connect with other TEX user associations around the world. Information about their activities is published regularly in TTN; reciprocal membership arrangements are available with some of the European groups.

7. TEX Training: Each year TUG offers seminars and classes for beginner, intermediate and advanced TEX users throughout North America and Europe. TUG will customize on-site courses to meet your company’s specific needs.

8. Discounts:

▷ Discounted Annual Meeting fees ▷ Student rates for membership (50% off individual rates!) ▷ Discounts on the purchase of selected TUG publications ▷ Discounts on TUG-sponsored courses

TEX Users Group  Phone: +1 (805) 963-1338
P.O. Box 869 FAX: +1 (805) 963-8358
Santa Barbara, CA 93102 USA E-mail: tug@math.ams.org